OpenStudy (anonymous):

Q'store = d(Qstore )/ dt = d(m·CC (T )·.Tr )/ dt where m is the mass of the conductor from which it is made the resistor, .Tr = Tr -T0 is the heating of the resistor, CC(T) is the specific heat (in function of the temperature) of the conductor, T0 is the initial temperature (considered equal with the temperature of the ambient medium). I search function Cp (T) for a nichrome resistor and thank you .

OpenStudy (anonymous):

Q'store = d(Qstore )/ dt = d(m·Cc (T )·.Tr )/ dt where m is the mass of the conductor from which it is made the resistor, .Tr = Tr -T0 is the heating of the resistor, Cc(T) is the specific heat (in function of the temperature) of the conductor, T0 is the initial temperature (considered equal with the temperature of the ambient medium). I search function Cc (T) for a nichrome resistor and thank you .